ITS Service Status Report

Service Degradation

Data Den Keyserver Issues.

Services Affected: Research Storage, Data Den Research Archive

Start Time: 05/18/2023 1:00 am

Anticipated End Time: Unknown

Issue Symptoms: Degradation

Update 5/24: We have taken three tape drives offline and one is due to replaced this afternoon or tomorrow.  Things are working better, if only in a performance degraded state with 3 of the 16 drives offline.  We are running some test pre-migrates to determine if this is viable until we can get the other drives replaced or diagnosed.

Update 5/23: We are still working with IBM to fix the issue.  Some large uploads may require temporary increases to cache to finish. Please let us know if you see any issues with longer uploads to Data Den.

Update 5/19:  We have restarted all GPFS services and removing some stuck snapsnots.   We're taking a look at things to make sure things are working as expected during the next round of data copying to tape. 

One of the two Data Den Keyservers is offline due to an as of yet unknown issue.   We are still diagnosing.  There should be no direct issues with accessing data as the other keyserver is still serving information.  

Who is Impacted? Users of the Data Den Research Archive

Next Update: 5/25 noon.

Technical Details

Service Type: Production

Report Additional Impacts

Contact the ITS Service Center for more information or to report additional impacts.